In the Bohr model, the electron of a hydrogen
atom moves in a circular orbit of radius
5.3 x 10^-11 m with a speed of 2.2 x 10^6 ms^-1. Determine its frequency and the current in the orbit.
1
Expert's answer
2019-05-06T10:29:44-0400

Frequency of the electron is:


"\nu=\frac{\omega}{2\pi}=\frac{v}{2\pi r}=\frac{2.2*10^{6}}{2*3.14 *5.3*10^{-11}}=6.6*10^{15} Hz"


The current in the orbit is:


"I=\frac{\Delta q}{\Delta t}=\frac{e}{T}=e\nu=1.6*10^{-19}*6.6*10^{15}=1.06*10^{-3} A"